# Mohammad Mehdi Morovati    Florian Tambon    Mina Taraghi    Amin Nikanjam    Foutse Khomh

### v2

@@ 2 Results
DRL Issues
Comprehension
48% 52% 69% 23% 8% 60% 23% 17%
Reward
86% 14% 32% 37% 31% 28% 28% 44%
Action
49% 51% 74% 11% 15% 66% 20% 14%
Environment
83% 17% 32% 34% 34% 31% 29% 40%
Hyperparameters
80% 20% 43% 32% 25% 37% 26% 37%
Design problem
75% 25% 38% 37% 25% 34% 31% 35%
Policy
62% 38% 54% 37% 9% 54% 32% 14%
State/Observation
68% 32% 52% 32% 16% 49% 37% 14%
DRL libraries/ frameworks
Installation
43% 57% 85% 12% 3% 78% 15% 7%
Dependency
54% 46% 78% 14% 8% 77% 18% 5%
API usage
38% 62% 86% 9% 5% 78% 14% 8%
Documentation
51% 49% 69% 23% 8% 71% 20% 9%
Bugs inside framework
48% 52% 75% 14% 11% 70% 15% 15%
Best fitted library
51% 49% 70% 25% 5% 65% 25% 10%
Parallel processing
GPU usage
54% 46% 72% 20% 8% 69% 20% 11%
Multi-threading
45% 55% 68% 24% 8% 65% 18% 17%
Multi-processing
49% 51% 57% 31% 12% 51% 23% 26%
Distributed systems
43% 57% 58% 28% 14% 55% 14% 31%
DL Issues
Model
48% 52% 71% 23% 6% 66% 26% 8%
Data prepossessing
48% 52% 57% 25% 18% 63% 25% 12%
DL framework
46% 54% 66% 25% 9% 64% 27% 9%
According to the survey results, the majority of our respondents have been confronted with challenges classified as DRL issues ($67.1\%$ average over all subcategories in this category). This observation is aligned with the proportion of SO posts categorized as DRL issues (see Fig.1). Conversely, DRL libraries/frameworks challenges have been experienced the least with only $46.7\%$ of respondents (which is the lowest proportion among all categories) reporting having faced challenges leveraging DRL libraries. This finding is reflected by the results of our quantitative analysis of SO posts, which show that only $21.6\%$ of posts contained questions related to the DRL libraries/frameworks category. It should be also taken into consideration that previous research showed that there is a growing trend toward the number of studies on DRL panzer2022deep ; kiran2021deep . Among the challenges identified in our taxonomy, developers who participated in our survey specify reward (86%), environment (83%), hyperparameters (80%), and design problem (75%) as the most common challenges in DRL development. Although only 14.8% of SO posts contained questions about API usage, 38% of survey respondents identified it as a challenging issue in DRL development. Given that reward, environment, hyperparameters, and design problem are fundamental components of an RL application lorenz2022reinforcement , it is expected that survey participants reported them as the most encountered challenges. For instance, defining the environment is known as a crucial step in an RL application development process that affects the convergence of an agent’s behavior significantly reda2020learning .
@@
We also ask survey respondents about the severity and needed effort to address the challenges identified in the provided taxonomy (Table 4). In general, the majority (exceeding $52\%$) of the survey respondents indicated that the most frequent challenges from their viewpoint (i.e., Reward, Environment, Design problem and Hyperparameters) are critical. Moreover, at least $59\%$ of the survey participants considered the level of effort required to address these challenges, to be “Medium” or “High”. The majority (more than $54\%$) of the survey participants consider the other identified challenges to be of Minor severity level, and to require a Low level of effort. In general, the participants consider that Installation and API usage challenges require a low level of effort, which might signal that DRL libraries/frameworks have good documentation and usability in general mojica2023machine .
@@
We extract information about the programming languages used to develop DRL applications from the collected SO posts. It should be mentioned that the posts were collected without any distinction on the programming language and frameworks used. Fig.5 presents the proportion of Python posts for the different identified categories of challenges. As can be seen, Python is by far the dominant programming language for all categories of challenges. However, the proportion of posts related to other programming languages and containing DRL issues is non-negligible (i.e., 20.2%). This high ratio is mostly attributable to Java (5.2% of all posts), C++ (4.7% of all posts), and R (4.7% of all posts) programming languages. It is also noteworthy that investigating the relationship between used programming languages and the challenges (subcategories) within each category reveals that Python stands out as the predominant programming language across all DRL development challenges. Based on these results, we conclude that there is no relationship between DRL development challenges and used programming languages. This finding is in accordance with prior research morovati2023bugs ; humbatova2020taxonomy which reported that Python is the most popular programming language for ML-enabled applications. These results about programming languages used in DRL applications development are also supported by our validation survey where all participants mentioned Python as the programming language they use for developing DRL applications. Besides, 20% of participants reported C/C++, and 12% mentioned other programming languages in addition to Python (e.g., C# 4%, and Java 3%).
@@
We examined the mentioned frameworks and libraries in posts related to different challenge categories. Fig.5(a) shows the number of times that each library/framework has been mentioned in posts belonging to various subcategories of the DRL issues category. With an exception for the environment and state/observation subcategories, Keras, Tensorflow, and PyTorch are the most used library/framework by SO users, and they are the most popular libraries/frameworks for developing ML and DLmorovati2024bug . Considering that there are several libraries/frameworks specifically designed to ease DRL application development (e.g. KerasRL, RLlib, etc), Fig.5(a) exposes that SO users usually prefer to use popular ML libraries/frameworks which can be leveraged to implement DRL applications as a subdomain of ML. On the other hand, gym brockman2016openai is the most popular library/framework in the environment subcategory which is reasonable as it is the most popular library for implementing various RL environments and provides a standard benchmark containing a large number of well-known RL environments panerati2021learning . This observation can be related to the inherent nature of SO posts, which may not necessarily provide details about the libraries/frameworks employed by the user posting questions. It is noteworthy that 46% of 885 examined SO posts lack any reference to the used DRL-related libraries/frameworks. As an example, post #56312962 serves as an illustrative case where no information has been mentioned regarding the utilized libraries-frameworks to implement the DRL application.
@@
Fig.5(b) presents the libraries/frameworks mentioned in the posts classified as DL issues. Results show that TensorFlow, Keras, and PyTorch are the most popular libraries/frameworks in the DL framework and model subcategories. Given that the challenges within the DL issues category pertain to the DL parts of DRL applications, it is not surprising to see TensorFlow, Keras, and PyTorch frequently mentioned since they are the most used libraries in the development of DL-enabled applications morovati2023bugs ; humbatova2020taxonomy . Moreover, the generality of Ray in the Data preprocessing subcategory, compared to other libraries/frameworks can be attributed to the fact that Ray encompasses not only DRL-related libraries (e.g., RLlib liang2018rllib ) but also various other libraries for a wide range of ML-related tasks at the same time, including scalable datasets, model training, and hyperparameter tuning pumperla2023learning which may ease developing DRL applications.
@@
About the most frequently referenced libraries/frameworks in the DRL libraries/frameworks category, as illustrated in Fig.5(c), gym received the largest number of questions, especially on the topics of API usage, installation, and dependency challenges. Although gym is the most popular library for implementing RL/DRL environments panerati2021learning , comparing Fig.5(a) and 5(b) may indicate that gym has less matured documentation and tutorials, in comparison to Keras, Tensorflow, and PyTorch. It is also worth mentioning that Keras, Tensorflow, and PyTorch are general-purpose ML-related libraries/frameworks which are more developed, compared to gym which is implemented specifically for RL development open_ai_gym .
@@ 3 Discussion
Regarding the provided taxonomy, it is worth mentioning that all of the categories and subcategories of challenges directly relate to the DRL application development. However, some of the challenges may be observed in other ML-related applications. For instance, all of the challenges belonging to DRL libraries/frameworks (e.g. API usage, Installation, Dependency, etc.) have been faced by all developers who use ML/DL libraries/frameworks. But it should be also taken into consideration that all of the investigated SO posts in this study have been achieved after a comprehensive filtering process making sure all of them are about challenges in DRL application development. On the other hand, Reward, Environment, Action, State/Observation, and Policy challenges are Specific to DRL applications. It should be also taken into consideration that what makes this taxonomy valuable in the DRL community is the fact that their frequency, importance, and severity would be different in DRL application development compared to other ML-related applications. For example, 14.8% of challenges in DRL application development are related to API usage, whereas it is only 5.3% in DL-related applications humbatova2020taxonomy .
In terms of the metrics employed to measure the difficulty of addressing challenges, our methodology aligns with prior studies analyzing SO posts alshangiti2019developing ; decan2019empirical . However, it is worth noting that some studies have utilized the number of posts within each category as a metric to illustrate the difficulty associated with addressing the related challenges bangash2019developers ; hamidi2021towards . In the case of considering the number of posts as an indicator of the difficulty level for tackling challenges (the number mentioned along with each category in Fig. 1), the results would closely mirror the difficulty levels observed for the challenges categorized as DL issues and parallel processing categories depicted in Fig.4. For instance, design problem is the second most difficult challenge in the DRL issues category when we use ‘number of posts’ or ‘duration to receive an accepted answer’ to measure the difficulty level. Conversely, the difficulty levels for challenges falling under DRL issues and DRL libraries/frameworks differ by using these two metrics. For example, as illustrated in Fig.3(b), hyperparameters emerges as the most challenging subcategory within the DRL issues category, even though comprehension has the highest number of posts. This discrepancy may be attributed to the nature of inquiring about hyperparameters, which may necessitate various implementations and a longer time to respond compared to other challenges. Furthermore, addressing SO posts categorized as comprehension mainly seeks users’ background knowledge, which does not necessarily require practical implementation or application execution.
@@
Finding 2 revealed that $27.3\%$ of DRL development challenges categorized as comprehension are related to the lack of sound understanding of basic DRL concepts. In other words, $58.4\%$ of posts belonging to DRL issues category (DRL-specific category) are about comprehension challenge. This finding highlights the need for documentation and tutorials to help DRL practitioners who are not experts in DRL, in the development of DRL applications. A roadmap for the development of DRL applications would also help developers navigate through the implementation of DRL applications with fewer misunderstandings of DRL concepts. The need for such material is emphasized by a post [^3] https://stackoverflow.com/questions/37973108 asking questions about the difference between RL and DRL. By providing a roadmap that systematically expands DRL developers’ understanding, developers will be supported in overcoming the most common challenge in DRL application development. An illustration of such guidance is the work conducted by Garg et al. garg2019roadmap on creating a roadmap for DL development. The need for good documentation and guidance is also emphasized by the survey participants who mentioned that ‘although there are a number of tutorials to start working on DRL, a few issues are shared between many of them’. The participants also noted that many of the DRL tutorial documents cover only a specific domain of DRL. Participants also lamented the poor usability of DRL-related tutorials, claiming that they often contain a lot of unnecessary materials.
@@
Our results in Finding 2 also stress the need for supporting tools and documentation, for parallel and distributed DRL application developments. Questions related to this topic took a long time before receiving an accepted answer on Stack Overflow. DRL experts could consider developing pre-configured packages to support parallel/distributed DRL application developments. In fact, Openja et al. openja2022studying examined ML application deployment practices on Docker and reported that a significant number of ML developers use Docker to manage dependencies, environment, and the execution of ML applications.
We have conducted a thorough examination of top DRL-related GitHub repositories to ensure the completeness of the generated taxonomy. Our methodology for selecting DRL-related repositories follows the established approach documented in similar prior studies morovati2023bugs ; morovati2024bug ; humbatova2020taxonomy . To this end, we use 7,244 repositories extracted to achieve contributors of DRL-related repositories (Subsection 1.2). From this pool, we select the top 100 repositories based on the highest number of stars. Upon reviewing the documentation and commit messages of these repositories, we do not observe any occurrences of mentioning challenges faced during the implementation of the DRL-related applications in these resources. Additionally, an exploration of raised issues in the selected repositories revealed a consistent pattern wherein users primarily seek support on the utilization of DRL applications offered by the repositories.
@@ 4.2 RL and DRL Quality Assurance
Nikanjam et al. nikanjam2022faults also investigated challenges categorized as DRL issues in our proposed taxonomy. They examined questions/discussions about four popular DRL frameworks (including gym brockman2016openai , Tensorforce tensorforce , Dopamine castro18dopamine , and Keras-rl plappert2016kerasrl on GitHub and SO) and extracted 329 SO posts about DRL. They categorized these posts into six groups: basic concepts, without acknowledgment, implementation issues, answered by the owner, relative questions, and others. They reported that ‘without acknowledgment’ and ‘implementation’ questions are the most common DRL-related questions in SO, accounting for 32% and 27%, respectively. They also showed that in 2% of their studied SO posts, the answer has been posted by the questioner. They report that DRL-related SO posts take an average and median time of 2.07 days and 13 hours, respectively, before receiving an accepted answer. This period is longer than the time taken by DL-related SO posts to receive an accepted answer; which is 5 hours on average. This finding implies that DRL-related questions might be more difficult to answer than DL-related questions. Nikanjam et al. nikanjam2022faults also proposed a taxonomy of faults in DRL models. with 11 different issue types. Our study differs from what Nikanjam et al. nikanjam2022faults carried out in several aspects. Firstly, they concentrated on four specific libraries/frameworks designed for DRL development, ignoring other SO posts containing source codes using other Python-based libraries/frameworks or not mentioning any script. For instance, a number of SO posts inquire about DRL concepts (e.g. #52838439) without mentioning any scripts, a scope not covered by Nikanjam et al. nikanjam2022faults . Moreover, their research was just on the DRL model, whereas our study delves into the challenges developers may face throughout the development of entire DRL applications, without any limitation to a specific section of DRL applications. Last but not least, Nikanjam et al. nikanjam2022faults examined SO posts reporting program faults during the development of DRL applications. It is crucial to note that challenges do not necessarily equate to program faults. As a result, we have included numerous SO posts in our studies that do not mention any program faults but rather pose questions related to understanding DRL concepts. Nikanjam et al. nikanjam2022faults also investigated challenges categorized as ‘DRL issues’ in our proposed taxonomy. They examined questions/discussions about four popular DRL frameworks (including gym brockman2016openai , Tensorforce tensorforce , Dopamine castro18dopamine , and Keras-rl plappert2016kerasrl on GitHub and SO) and extracted 329 SO posts about DRL. They categorized these posts into six groups: basic concepts, without acknowledgment, implementation issues, answered by the owner, relative questions, and others. They reported that ‘without acknowledgment’ and ‘implementation’ questions are the most common DRL-related questions in SO, accounting for $32\%$ and $27\%$, respectively. They also showed that in $2\%$ of their studied SO posts, the answer has been posted by the questioner. They report that DRL-related SO posts take an average and median time of $2.07$ days and $13$ hours, respectively, before receiving an accepted answer. This time period is longer than the time taken by DL-related SO posts to receive an accepted answer; which is $5$ hours on average. This finding implies that DRL-related questions might be more difficult to answer than DL-related questions. Nikanjam et al. nikanjam2022faults also proposed a taxonomy of faults in DRL models. The most significant difference between their study and ours is their focus on the DRL model only and the fact that they only collected data about faults mentioned in SO; disregarding all the other types of questions.

### v3

@@ 2 Results
DRL Issues Comprehension 48% 52% 69% 23% 8% 60% 23% 17%
Reward 86% 14% 32% 37% 31% 28% 28% 44%
Action 49% 51% 74% 11% 15% 66% 20% 14%
Environment 83% 17% 32% 34% 34% 31% 29% 40%
Hyperparameters 80% 20% 43% 32% 25% 37% 26% 37%
Design problem 75% 25% 38% 37% 25% 34% 31% 35%
Policy 62% 38% 54% 37% 9% 54% 32% 14%
State/Observation 68% 32% 52% 32% 16% 49% 37% 14%
DRL libraries/ frameworks Installation 43% 57% 85% 12% 3% 78% 15% 7%
Dependency 54% 46% 78% 14% 8% 77% 18% 5%
API usage 38% 62% 86% 9% 5% 78% 14% 8%
Documentation 51% 49% 69% 23% 8% 71% 20% 9%
Bugs inside framework 48% 52% 75% 14% 11% 70% 15% 15%
Best fitted library 51% 49% 70% 25% 5% 65% 25% 10%
Parallel processing GPU usage 54% 46% 72% 20% 8% 69% 20% 11%
Multi-threading 45% 55% 68% 24% 8% 65% 18% 17%
Multi-processing 49% 51% 57% 31% 12% 51% 23% 26%
Distributed systems 43% 57% 58% 28% 14% 55% 14% 31%
DL Issues Model 48% 52% 71% 23% 6% 66% 26% 8%
Data prepossessing 48% 52% 57% 25% 18% 63% 25% 12%
DL framework 46% 54% 66% 25% 9% 64% 27% 9%
According to the survey results, the majority of our respondents have been confronted with challenges classified as DRL issues ($68.9\%$ average over all subcategories in this category). This observation is aligned with the proportion of SO posts categorized as DRL issues (see Fig.1). Conversely, challenges belong to Parallel Processing & Multi-threading category have been experienced the least with only $45.25\%$ of respondents (which is the lowest proportion among all categories) reporting having faced challenges leveraging parallel processing. This finding is reflected by the results of our quantitative analysis of SO posts, which show that only $1.5\%$ of posts contained questions related to the Parallel Processing & Multi-threading category. It should be also taken into consideration that previous research showed that there is a growing trend toward the number of studies on DRL panzer2022deep ; kiran2021deep . Among the challenges identified in our taxonomy, developers who participated in our survey specify reward (86%), environment (83%), hyperparameters (80%), and design problem (75%) as the most common challenges in DRL development. Although only 14.8% of SO posts contained questions about API usage, 38% of survey respondents identified it as a challenging issue in DRL development. Given that reward, environment, hyperparameters, and design problem are fundamental components of an RL application lorenz2022reinforcement , it is expected that survey participants reported them as the most encountered challenges. For instance, defining the environment is known as a crucial step in an RL application development process that affects the convergence of an agent’s behavior significantly reda2020learning .
@@
To enhance the completeness of our provided taxonomy, we scrutinize 254 sampled commits and issues extracted from real-world DRL-related repositories (Subsection 1.2). Upon thorough examination of the sampled commits, we did not identify any instances of challenges being mentioned in relation to the DRL. Furthermore, an analysis of sampled closed issues revealed a consistent pattern wherein users primarily seek support on the utilization of DRL applications offered by the repositories.
@@
We also ask survey respondents about the severity and needed effort to address the challenges identified in the provided taxonomy (Table 4). In general, the majority (exceeding $57\%$) of the survey respondents indicated that the most frequent challenges from their viewpoint (i.e., Reward, Environment, Design problem and Hyperparameters) are major or critical. Moreover, at least $63\%$ of the survey participants considered the level of effort required to address these challenges, to be “Medium” or “High”. The majority (more than $52\%$) of the survey participants consider the other identified challenges to be of Minor severity level, and to require a Low level of effort. In general, the participants consider that Installation and API usage challenges require a low level of effort, which might signal that DRL libraries/frameworks have good documentation and usability in general mojica2023machine .
@@
We extract information about the programming languages used to develop DRL applications from the collected SO posts. It should be mentioned that the posts have been collected without any distinction on the programming language and frameworks used. Fig.5 presents the proportion of posts using Python programming language for the different identified categories of challenges. As can be seen, Python is by far the dominant programming language for all categories of challenges. However, the proportion of posts mentioning other programming languages and containing DRL issues is non-negligible (i.e., 20.2%). This high ratio is mostly attributable to Java (5.2% of all posts), C++ (4.7% of all posts), and R (4.7% of all posts) programming languages. It is also noteworthy that investigating the relationship between used programming languages and the challenges (subcategories) within each category reveals that Python stands out as the predominant programming language across all DRL development challenges. Based on these results, we conclude that there is no relationship between DRL development challenges and used programming languages. This finding is in accordance with prior research morovati2023bugs ; humbatova2020taxonomy which reported that Python is the most popular programming language for ML-enabled applications. These results about programming languages used in DRL applications development are also supported by our validation survey where all participants mentioned Python as the programming language they use for developing DRL applications. Besides, 20% of participants reported C/C++, and 12% mentioned other programming languages in addition to Python (e.g., C# 4%, and Java 3%).
@@
We also examine the mentioned frameworks and libraries in posts related to different challenge categories. Fig.6(a) shows the number of times that each library/framework has been mentioned in posts belonging to various subcategories of the DRL issues category. With an exception for the environment and state/observation subcategories, Keras, Tensorflow, and PyTorch are the most used libraries/frameworks by SO users, and they are the most popular libraries/frameworks for developing ML and DLmorovati2024bug . Considering that there are several libraries/frameworks specifically designed to ease DRL application development (e.g. KerasRL, RLlib, etc), Fig.6(a) exposes that SO users usually prefer to use popular ML libraries/frameworks which can be leveraged to implement DRL applications as a subdomain of ML. On the other hand, gym brockman2016openai is the most popular library/framework in the environment subcategory which is reasonable as it is the most popular library for implementing various RL environments and provides a standard benchmark containing a large number of well-known RL environments panerati2021learning . This observation can be related to the inherent nature of SO posts, which may not necessarily provide details about the libraries/frameworks employed by the user posting questions. It is noteworthy that 46% of 885 examined SO posts lack any reference to the used DRL-related libraries/frameworks. As an example, post #56312962 serves as an illustrative case where no information has been mentioned regarding the utilized libraries-frameworks to implement the DRL application.
@@
Fig.6(b) presents the libraries/frameworks mentioned in the posts classified as DL issues. Results show that TensorFlow, Keras, and PyTorch are the most popular libraries/frameworks in the DL framework and model subcategories. Given that the challenges within the DL issues category pertain to the DL parts of DRL applications, it is not surprising to see TensorFlow, Keras, and PyTorch are frequently mentioned since they are the most used libraries in the development of DL-enabled applications morovati2023bugs ; humbatova2020taxonomy . Moreover, the generality of Ray in the Data preprocessing subcategory, compared to other libraries/frameworks can be attributed to the fact that Ray encompasses not only DRL-related libraries (e.g., RLlib liang2018rllib ) but also various other libraries for a wide range of ML-related tasks at the same time, including scalable datasets, model training, and hyperparameter tuning pumperla2023learning which may ease developing DRL applications.
@@
About the most frequently referenced libraries/frameworks in the DRL libraries/frameworks category, as illustrated in Fig.6(c), gym received the largest number of questions, especially on the topics of API usage, installation, and dependency challenges. Although gym is the most popular library for implementing RL/DRL environments panerati2021learning , comparing Fig.6(a) and 6(b) may indicate that gym has less matured documentation and tutorials, in comparison to Keras, Tensorflow, and PyTorch. It is also worth mentioning that Keras, Tensorflow, and PyTorch are general-purpose ML-related libraries/frameworks which are more developed, compared to gym which is implemented specifically for RL development open_ai_gym .
@@ 3 Discussion
Regarding the provided taxonomy, it is worth mentioning that all of the categories and subcategories of challenges directly relate to the DRL application development. However, some of the challenges may be observed in other ML-related applications. For instance, all of the challenges belonging to DRL libraries/frameworks (e.g. API usage, Installation, Dependency, etc.) have been faced by all developers who use ML/DL libraries/frameworks. But it should be also taken into consideration that all of the investigated SO posts in this study have been achieved after a comprehensive filtering process making sure all of the extracted SO posts are about challenges in DRL application development. On the other hand, Reward, Environment, Action, State/Observation, and Policy challenges are Specific to DRL applications. It should be also taken into consideration that what makes this taxonomy valuable in the DRL community is the fact that their frequency, importance, and severity would be different in DRL application development compared to other ML-related applications. For example, 14.8% of challenges in DRL application development are related to API usage, whereas it is only 5.3% in DL-related applications humbatova2020taxonomy .
@@
Finding 2 revealed that $27.3\%$ of DRL development challenges categorized as comprehension are related to the lack of sound understanding of basic DRL concepts. In other words, $58.4\%$ of posts belonging to DRL issues category (DRL-specific category) are about comprehension challenge. This finding highlights the need for documentation and tutorials to help DRL developers who are not experts in DRL, in the development of DRL applications. A roadmap for the development of DRL applications would also help developers navigate through the implementation of DRL applications with fewer misunderstandings of DRL concepts. The need for such material is emphasized by a post [^3] https://stackoverflow.com/questions/37973108 asking questions about the difference between RL and DRL. By providing a roadmap that systematically expands DRL developers’ understanding, developers will be supported in overcoming the most common challenge in DRL application development. An illustration of such guidance is the work conducted by Garg et al. garg2019roadmap on creating a roadmap for DL development. The need for good documentation and guidance is also emphasized by the survey participants who mentioned that ‘although there are a number of tutorials to start working on DRL, a few issues are shared between many of them’. The participants also noted that many of the DRL tutorial documents cover only a specific domain of DRL. Participants also lamented the poor usability of DRL-related tutorials, claiming that they often contain a lot of unnecessary materials.
@@
Our results in Finding 2 also stress the need for supporting tools and documentation, for parallel and distributed DRL application developments. Questions related to this topic took a long time before receiving an accepted answer on Stack Overflow. DRL experts could consider developing pre-configured packages to support parallel/distributed DRL application developments. As an example of the same task in the ML-related development, Openja et al. openja2022studying examined ML application deployment practices on Docker and reported that a significant number of ML developers use Docker to manage dependencies, environment, and the execution of ML applications.
@@ 4.2 RL and DRL Quality Assurance
Nikanjam et al. nikanjam2022faults also investigated challenges categorized as DRL issues in our proposed taxonomy. They examined questions/discussions about four popular DRL frameworks (including gym brockman2016openai , Tensorforce tensorforce , Dopamine castro18dopamine , and Keras-rl plappert2016kerasrl on GitHub and SO) and extracted 329 SO posts about DRL. They categorized these posts into six groups: basic concepts, without acknowledgment, implementation issues, answered by the owner, relative questions, and others. They reported that ‘without acknowledgment’ and ‘implementation’ questions are the most common DRL-related questions in SO, accounting for 32% and 27%, respectively. They also showed that in 2% of their studied SO posts, the answer has been posted by the questioner. They report that DRL-related SO posts take an average and median time of 2.07 days and 13 hours, respectively, before receiving an accepted answer. This period is longer than the time taken by DL-related SO posts to receive an accepted answer; which is 5 hours on average. This finding implies that DRL-related questions might be more difficult to answer than DL-related questions. Nikanjam et al. nikanjam2022faults also proposed a taxonomy of faults in DRL models. with 11 different issue types. Our study differs from what Nikanjam et al. nikanjam2022faults carried out in several aspects. Firstly, they concentrated on four specific libraries/frameworks designed for DRL development, ignoring other SO posts containing source codes using other Python-based libraries/frameworks or not mentioning any script. For instance, a number of SO posts inquire about DRL concepts (e.g. #52838439) without mentioning any scripts, a scope not covered by Nikanjam et al. nikanjam2022faults . Moreover, their research was just on the DRL model, whereas our study delves into the challenges developers may face throughout the development of entire DRL applications, without any limitation to a specific section of DRL applications. Last but not least, Nikanjam et al. nikanjam2022faults examined SO posts reporting program faults during the development of DRL applications. It is imperative to distinguish between challenges and program faults and note that challenges do not necessarily equate to program faults. A fault denotes a defect or error leading to a discrepancy between the expected and achieved results or observed behavior morovati2023bugs . Software development challenges, in contrast, encompass any difficulty or complexity encountered in completing a development task. These challenges may arise from various factors, including technical complexities, resource constraints, or lack of expertise. Nikanjam et al. nikanjam2022faults also investigated challenges categorized as ‘DRL issues’ in our proposed taxonomy. They examined questions/discussions about four popular DRL frameworks (including gym brockman2016openai , Tensorforce tensorforce , Dopamine castro18dopamine , and Keras-rl plappert2016kerasrl on GitHub and SO) and extracted 329 SO posts about DRL. Since Nikanjam et al. nikanjam2022faults focused on software faults, they excluded 305 SO posts from their dataset, generating their taxonomy based on 24 such posts that included faults explicitly. Notably, all SO posts in their dataset are included in our study. Furthermore, our research incorporates SO posts that do not pertain to program faults but instead pose queries related to the comprehension of DRL concepts. For example, the comprehension subcategory includes 253 SO posts, none of which reports program faults. They categorized these posts into six groups: basic concepts, without acknowledgment, implementation issues, answered by the owner, relative questions, and others. They reported that ‘without acknowledgment’ and ‘implementation’ questions are the most common DRL-related questions in SO, accounting for $32\%$ and $27\%$, respectively. They also showed that in $2\%$ of their studied SO posts, the answer has been posted by the questioner. They report that DRL-related SO posts take an average and median time of $2.07$ days and $13$ hours, respectively, before receiving an accepted answer. This time period is longer than the time taken by DL-related SO posts to receive an accepted answer; which is $5$ hours on average. This finding implies that DRL-related questions might be more difficult to answer than DL-related questions. Nikanjam et al. nikanjam2022faults also proposed a taxonomy of faults in DRL models. The most significant difference between their study and ours is their focus on the DRL model only and the fact that they only collected data about faults mentioned in SO; disregarding all the other types of questions.